Five cubic feet of an aqueous hydrochloric acid (H
Answer #1

1 cubis feet = 28.3 Lit or 28.3 kg of water

5 cubic feet = 5 X 28.3 = 141.5 lit of water having 2M of HCl

HCl moleculer weight + 36.46

Molarity = W/M.WtX 1/Lit

2 = W/ 36.46 X 1/141.5

a). Weight of HCl = 2 X 36.46 X141.5 = 10318.18gr

b). 141.5 lit of water need

c). mole fraction of HCl

HCL moles = 10318.18/36.46 = 282.99 mole

H2O moles = 141500/18 = 7861.11 moles

mole fraction of HCl = 282.99/ 282.99+7861.11= 0.035